# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

### v2

@@ IV Summary and outlook
Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [33, 17] or charge-state classifications [34, 24]. However, these algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed. BATIS fills this gap, bringing device characterization easily within reach for linear QD arrays and paving the way toward fully automatic control. When linked together, the proposed subroutines form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, putting QD qubit systems on the path to scalable quantum computing.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work was supported in part by ARO grant nos. W911NF-17-1-0274, W911NF-23-1-0110, and W911NF-24-2-0043. The authors gratefully acknowledge the use of facilities and instrumentation in the UW-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF DMR-2309000) and the University of Wisconsin-Madison. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Reviews of Modern Physics 95, 025003 (2023).
@@
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, et. al, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022), pMID: 34806894.
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, et. al, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, et. al, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
- Note [2] For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable. However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.
- Borsoi et al. [2024] F. Borsoi, N. W. Hendrickx, V. John, M. Meyer, S. Motz, F. Van Riggelen, A. Sammak, S. L. De Snoo, G. Scappucci, and M. Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nature Nanotechnology 19, 21 (2024).
- Park et al. [2024] S. Park, J. Benson, J. Corrigan, J. P. Dodson, S. N. Coppersmith, M. Friesen, and M. A. Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing (2024), arXiv:2408.15380 [cond-mat].
@@
- de Kruijf et al. [2023] M. de Kruijf, S. Geyer, T. Berger, M. Mergenthaler, F. Braakman, R. J. Warburton, and A. V. Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Review of Scientific Instruments 94, 054707 (2023).
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, et. al, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
- Zubchenko et al. [2024] A. Zubchenko, D. Middlebrooks, T. Rasmussen, L. Lausen, F. Kuemmeth, A. Chatterjee, and J. P. Zwolak, Autonomous bootstrapping of quantum dot devices, arXiv:2407.20061 (2024).
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, et. al, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
- Wolfe et al. [2024] M. A. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. G. Lagally, R. McDermott, M. Friesen, et. al, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
- Rao et al. [2024] A. S. Rao, D. Buterakos, B. van Straaten, V. John, C. X. Yu, S. D. Oosterhout, L. Stehouwer, G. Scappucci, M. Veldhorst, F. Borsoi, and J. P. Zwolak, MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays, arXiv:2411.12516 (2024).
@@
- Schug et al. [2024] D. Schug, T. J. Kovach, M. A. Wolfe, J. Benson, S. Park, J. P. Dodson, J. Corrigan, M. A. Eriksson, and J. P. Zwolak, Automation of Quantum Dot Measurement Analysis via Explainable Machine Learning (2024), version Number: 4.

### v3

@@ V Summary and outlook
Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [47, 22] or charge-state classifications [14, 16]. These algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed. BATIS helps to fill this gap, enabling bootstrapping of gate-defined devices in which current paths are entirely defined by gate voltages whose values cannot be known in advance.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC, for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work was supported in part by ARO grant nos. W911NF-17-1-0274, W911NF-23-1-0110, and W911NF-24-2-0043. The authors gratefully acknowledge the use of facilities and instrumentation in the UW-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF DMR-2309000) and the University of Wisconsin-Madison. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Rev. Mod. Phys. 95, 025003 (2023).
@@
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, et al., A flexible design platform for Si/SiGe exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022) .
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, et al., Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, et al., 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line, arXiv:2410.16583 (2024) .
- Zwolak et al. [2024] J. P. Zwolak, J. M. Taylor, R. W. Andrews, J. Benson, G. W. Bryant, D. Buterakos, A. Chatterjee, S. Das Sarma, M. Eriksson, E. Greplová, et al., Data needs and challenges for quantum dot devices automation, npj Quantum Inf. 10, 105 (2024).
- Note [2] Mathematically, the pinch-off and the turn-on voltage are indistinguishable for a given current curve. However, when measuring a physical gate, a hysteresis can be observed.
- Rao et al. [2024] A. S. Rao, D. Buterakos, B. van Straaten, V. John, C. X. Yu, S. D. Oosterhout, L. Stehouwer, G. Scappucci, M. Veldhorst, F. Borsoi, and J. P. Zwolak, Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Array, Phys. Rev. X 15, 021034 (2025).
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, et al., Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
- Zubchenko et al. [2024] A. Zubchenko, D. Middlebrooks, T. Rasmussen, L. Lausen, F. Kuemmeth, A. Chatterjee, and J. P. Zwolak, Autonomous bootstrapping of quantum dot devices, Phys. Rev. Appl. 23, 014072 (2025).
- Borsoi et al. [2023] F. Borsoi, N. W. Hendrickx, V. John, M. Meyer, S. Motz, F. van Riggelen, A. Sammak, S. L. de Snoo, G. Scappucci, and M. Veldhorst, Shared control of a 16 semiconductor quantum dot crossbar array, Nat. Nanotechnol. 19, 21 (2023).
- Park et al. [2024] S. Park, J. Benson, J. Corrigan, J. P. Dodson, S. N. Coppersmith, M. Friesen, and M. A. Eriksson, Single-shot latched readout of a quantum dot qubit using barrier gate pulsing, arXiv:2408.15380 (2024).
@@
- de Kruijf et al. [2023] M. de Kruijf, S. Geyer, T. Berger, M. Mergenthaler, F. Braakman, R. J. Warburton, and A. V. Kuhlmann, A compact and versatile cryogenic probe station for quantum device testing, Rev. Sci. Instrum. 94, 054707 (2023).
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, et al., Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
- Paurevic et al. [2025] A. Paurevic, A. Sakr, T. Joshi, D. van der Bovenkamp, Q. T. Nicolau, F. A. Zwanenburg, and J. Baugh, Automated tuning and characterization of a single-electron transistor charge sensor, arXiv:2502.02521 (2025) .
- Sze and Ng [2006] S. M. Sze and K. K. Ng, Physics of Semiconductor Devices (Wiley, New York, 2006).
- Massai et al. [2024] L. Massai, B. Hetényi, M. Mergenthaler, F. J. Schupp, L. Sommer, S. Paredes, S. W. Bedell, P. Harvey-Collard, G. Salis, A. Fuhrer, and N. W. Hendrickx, Impact of interface traps on charge noise and low-density transport properties in Ge/SiGe heterostructures, Commun. Mater. 5, 151 (2024).
- Rudolph et al. [2019] M. Rudolph, B. Sarabi, R. Murray, M. S. Carroll, and N. M. Zimmerman, Long-term drift of Si-MOS quantum dots with intentional donor implants, Sci. Rep. 9, 7656 (2019).
- Meyer et al. [2023] M. Meyer, C. Déprez, I. N. Meijer, F. K. Unseld, S. Karwal, A. Sammak, G. Scappucci, L. M. K. Vandersypen, and M. Veldhorst, Single-electron occupation in quantum dot arrays at selectable plunger gate voltage, Nano Lett. 23, 11593 (2023) .
- Wolfe et al. [2024a] M. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. Lagally, R. McDermott, M. Friesen, et al., Control of threshold voltages in /0.70.3 quantum devices via optical illumination, Phys. Rev. Appl. 22, 034044 (2024a).
@@
- [37] See Supplemental Material at URL-will-be-inserted-by-publisher.
- Wang et al. [2022] K. Wang, G. Xu, F. Gao, H. Liu, R.-L. Ma, X. Zhang, Z. Wang, G. Cao, T. Wang, J.-J. Zhang, et al., Ultrafast coherent control of a hole spin qubit in a germanium quantum dot, Nat. Commun. 13, 206 (2022).
- Tormo-Queralt et al. [2022] R. Tormo-Queralt, C. B. Møller, D. A. Czaplewski, G. Gruber, M. Cagetti, S. Forstner, N. Urgell-Ollé, J. A. Sanchez-Naranjo, C. Samanta, C. S. Miller, and A. Bachtold, Novel nanotube multiquantum dot devices, Nano Lett. 22, 8541 (2022) .
- Dirac and Fowler [1926] P. A. M. Dirac and R. H. Fowler, On the theory of quantum mechanics, Proc. R. Soc. Lond. Ser. A-Contain. Pap. Math. Phys. Character. 112, 661 (1926).
- Zannoni [1999] A. Zannoni, On the quantization of the monoatomic ideal gas, arXiv:cond-mat/9912229 (1999).
- Bardeen [1947] J. Bardeen, Surface states and rectification at a metal semi-conductor contact, Phys. Rev. 71, 717 (1947).
- Dacey and Ross [1953] G. C. Dacey and I. M. Ross, Unipolar "field-effect" transistor, Proc. IRE 41, 970 (1953).
@@
- Schug et al. [2024] D. Schug, T. J. Kovach, M. A. Wolfe, J. Benson, S. Park, J. P. Dodson, J. Corrigan, M. A. Eriksson, and J. P. Zwolak, Automation of quantum dot measurement analysis via explainable machine learning, Mach. Learn.: Sci. Technol. 6, 015006 (2025).